OpenStudy (anonymous):

Solve the equation by graphing:x^2+7x+19=0 a.x=49 b.no solution c.x=19 x=12

OpenStudy (anonymous):

y=x^2+7x+19 <-- graph that equation The solutions are where x=0 AKA where the line intersects the x axis

OpenStudy (anonymous):

You can find the solution by factoring or using the quadratic formula \[x = -b \pm \sqrt{b ^{2}-4ac}\] all over 2a

OpenStudy (anonymous):

I cant see wher it crosses 0 on my calculator

OpenStudy (anonymous):

try this with the quad formula: -7 +- √7^2 - 4(1)(19) /2

OpenStudy (anonymous):

7^2 = 49 4(1)(19) = 76 49 - 76 = -27 you cant have a negative under the radical so there are NO SOLUTIONS! Hope I helped!!

OpenStudy (anonymous):

By the way what I did above was the radical part of the formula...just by knowing that it is a negative will let you know that there are no solutions for the equation. Trust your calculator on solutions:) it was right, the equation doesnt cross the axis

OpenStudy (anonymous):

so the answer is B right @jazy @timo86m